OpenStudy (anonymous):

Write three ratios equivelent to 7/35

OpenStudy (anonymous):

@iambatman @perl @flasht @ilovewolf @haleyelizabeth2017 @One098 @jabez177 @Maria195 @JoeJoldin @partygirl @mathmate @Maria195

OpenStudy (anonymous):

@HelpBlahBlahBlah @perl @partygirl @mathmate @flasht @AnonymousBella @Evil-cow

OpenStudy (anonymous):

You can divide both sides by a common factor, or multiply both sides by the same number. They will both give you equal ratios. So for example 8/48 can also be written as 1/6 if both sides are divided by 8. It could also be multiplied to make 16/96 by multiplying by 2. Hopefully this helps :)

OpenStudy (mathmate):

Example: \(\Large \frac{1}{2}\equiv \frac{3}{6} \equiv \frac{5}{10}\) @kittenlover731 rememberif you tag a lot of people at any one time, people will delegate the help to the others, and you are likely to get less people to help you.